OpenStudy (anonymous):

A 2.00L glass soda bottle filled only with air is tightly capped at 24 degrees C and 744.0 mmHg. If the bottle is placed in water at 75 degrees C what is the pressure of the bottle

OpenStudy (cuanchi):

P1= 744.0 mmHg T1= 24C +273 = 297 K T2= 75C + 273= 348 K P2= ? P1/T1 = P2/T2 isolate P2 and solve the equation with the values of temperate in K your answer will be in mmHg